The test statistics of z= 2.21 is obtained when testing the claim that p>0.3

a. Identify the hypothesis test as being two tailed, left tailed, or right tailed

b. Find p value

c.using a significant level of = 0.05, should we reject H0 or should we fail To reject H0?
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.3

Ha : p > 0.3

z=2.21

P(z >2.21 ) = 1 - P(z < 2.21) = 1-0.9864=0.0136

P-value = 0.0136

= 0.05

P-value <

Reject the null hypothesis .